What is the density of oxygen o2 in grams per liter at 25°c and 0.85 atm


1
Expert's answer
2021-01-11T03:48:52-0500

Assuming you have 1mole of oxygen.

The volume of 1 mole of oxygen at stp = 22.4L

Mass of 1 mole of oxygen = 16g


stp means that temperature is 25oC and pressure is 1 atm.

In this case, the temperature remains the same (25oC) but pressure reduces.


We need to find the new volume of the gas at the reduced pressure.


We use Boyles Law;

P1V1 = P2V2

P1= 1 atm

P2 = 0.85 atm

V1 = 22.4L

V2 =?

V2 = "\dfrac{P1V1}{P2} = \dfrac{1atm x 22.4L}{0.85atm} = 26.353L"


Density = "\dfrac{mass}{volume} = \dfrac{16g}{26.353L} = 0.607g/L"